the problem is: 2y^2=14y. I need to solve this by factoring. The first thing I think I would do is factor out the greatest common factor which would be 2. I get to the point of 2(y^2-7y) but I am stuck on what y equals.
thanks,
Michael

Answer by Nate(3500)

2y^2 = 14y
2y^2 - 14y = 0
2y(y - 7) = 0
Anything times zero is equal to zero.
2y = 0
or y = 0
~ and ~
y - 7 = 0
or y = 7